# Enhancing Hokkien Dual Translation by Exploring and Standardizing of Four Writing Systems

### v1

@@ 5.1. Experimental Ablation Studies
In order to isolate the impact of various data inputs, we conducted ablation studies on continued pre-training of three different Hokkien writing systems, the extension of the input vocabulary with Hokkien Chinese characters and Hokkien Latin systems, and fine-tuning these three different Hokkien writing systems111111Our focus on Hokkien LRL led us to exclude HRL data from training and evaluation..
@@ 5.1.1. Continued Pre-training Corpora and Vocabulary Extension Ablation Studies
As shown in Table 5, we took LLaMA 2-7B without any continued pre-training as a baseline. We compared its performance with the TAIDE-7B model continued pre-trained on different monolingual data and evaluated the impact of extending the Hokkien dictionary. Here, “NONE” indicates no continued pre-training on any Hokkien data, “HAN” indicates pre-training solely on Chinese character-based Hokkien data, and “ALL” indicates the inclusion of Latin script (POJ) and hybrid (HL) Hokkien data in addition to “HAN”.
The baseline model, which has no further continued pre-training, performs the worst. In contrast, the TAIDE-7B model improves significantly in all translation directions without relying on Hokkien monolingual data but ZH data. This suggests that using a similar HRL model as a foundational model is beneficial when supplementary monolingual data is not available. Pre-training on HAN data improves the GPT-4 score by 4 to 6 points in HAN-related translations. Incorporating all Hokkien data yields the best performance, particularly in POJ-related translations, with a 10 to 20 point increase in GPT-4 score. These findings align with previous research (Yang et al., 2023; Li et al., 2023; Xu et al., 2023), demonstrating the substantial performance improvement in translation tasks when monolingual data is utilized for languages that the foundational model is not familiar with.
Regarding vocabulary extension, since the added vocabulary primarily consists of Latin scripts from POJ, the model with vocabulary extension exhibit superior performance only when the target language is POJ. For other translation directions, these models exhibit a slight decrement, averaging 3 points lower on the GPT-4 score compared to models without vocabulary extension. We attribute this to the limited size of the pre-training corpus, which hinders effective tuning of newly added tokens. Consequently, we opted not to extend the vocabulary and suggest future work in collecting a larger POJ corpus for further investigation.
@@ 5.1.2. Fine-tuning Datasets Ablation Study
Table 6 indicates that models fine-tuned on parallel datasets significantly outperform few-shot ICL models across all translation directions, particularly in directions involving Latin script Hokkien (POJ), with a GPT-4 score increase from 17 to 54 points. Demonstrate that the benefits of fine-tuning are particularly substantial for writing systems that are not closely related to HRL.
For ZHHAN and ENHAN directions, including ENHAN data enhance translation performance. However, further incorporation of POJ and HL parallel data does not yield additional improvements, suggesting that focusing solely on HRL in the parallel data is more effective for aligning cross-lingual embeddings.
In Hokkien script translation, the model performs the best when all parallel datas are included. The inclusion of POJHL data dramatically improve translation capabilities in the POJHAN. The chrF++ improved by 6.05 and 17.16 points for HAN-POJ and POJ-HAN directions, respectively. We hypothesize that HL contains a few Han characters, allowing the model to learn some lexical correlations between different scripts in Hokkien.
@@ 5.2. Pre-training Corpus Script-Standardization
We evaluated the efficacy of continued pre-training using three different datasets: 1) monolingual Hokkien data in Chinese characters only, 2) monolingual Hokkien data in all writing systems, and 3) all monolingual Hokkien data standardized into Han characters. These three models were subsequently fine-tuned using only ZHHAN and ENHAN parallel data.
Table 7 present the evaluation results of these three models on TAT. The model pre-trained on standardized monolingual data on par or slightly outperforms the other models on average across both pairs of languages. This suggests that standardizing the writing system into Chinese character can yield benefits on HAN translation.
@@ 6. Conclusion
We conducted a thorough examination of the efficiency of large language models in a translation system including ZH, English and two different Hokkien writing systems (HAN and POJ). We employed evaluation metrics including BLEU, chrF++, and a GPT-4-based scoring system. Our results showed that employing Hokkien similar high-resource language model as a foundational model led to significant performance enhancements. Furthermore, models pre-trained on all available Hokkien data exhibited the highest performance. Extending vocabulary appeared to favor the incorporation of more monolingual data to elicit their efficiency. In the fine-tuning stage, the model demonstrates enhanced HAN translation only when it is supplied with parallel data corresponding to the HRLs that were extensively involved in its prior pre-training. Additionally, we investigated the benefits of standardizing the Hokkien scripts to Chinese characters. Our future work will explore the potential advantage of data augmentation through translations from ZH monolingual corpora into Hokkien Han and assess its impact on translation quality. Moreover, extending this research to include other prevalent spoken languages in Taiwan, like Hakka, can offer a more extensive viewpoint on handling the linguistic variations in Taiwan.
@@ 7. Limitation
The methodology used in this study is conditional on the fact that one writing system in a low-resource language is similar to a high-resource language, which can be seen as a limitation. In this work, Hokkien Han and Mandarin Chinese share similar writing systems and possess a substantial amount of common vocabulary. This allows for a transfer of knowledge from extensive Mandarin Chinese texts, thereby leveraging the benefits of large language models pre-trained on abundant Mandarin Chinese corpora to achieve an exceptionally efficient translation system.
@@ 9. Acknowledgements
We would like to express our sincere gratitude to Dr. Yu-Chun Wang for the generous assistance and insightful discussions on Taiwanese Hokkien. Our thanks also go to Ray Hung for his invaluable help in polishing the paper writing. Additionally, we are deeply grateful to all the reviewers for their dedication to improving this research. We are thankful to the Trustworthy AI Dialogue Engine (TAIDE) project for providing the Traditional Chinese foundation LLM, which served as a crucial base for building our model. Special appreciation goes to the National Center for High-performance Computing (NCHC) for providing computational and storage resources. This work was also supported in part by the Co-creation Platform of the Speech-AI Research Center, Industry-Academia Innovation School, NYCU, under the framework of the National Key Fields Industry-University Cooperation and Skilled Personnel Training Act, from the Ministry of Education (MOE), the National Development Fund (NDF), and industry partners in Taiwan.
@@ 10. Bibliographical References
[…]
@@
Appendix A JSD for Corpora
A.1. JSD of Continued Pre-training Monolingual Data
We use Jensen-Shannon divergence (JSD) as a metric to assess the domain similarity of the HAN dataset. Since there are currently no word segmentation tools available for Hokkien writing systems, the calculation of JSD is a challenging task. The HAN dataset has a certain degree of similarity with Traditional Chinese characters. Therefore, we use the CKIP Traditional Chinese segmentation tool121212https://github.com/ckiplab/ckiptagger to segment the HAN dataset and compute the JSD to analyse the domain similarity of our dataset.
Figure 3 demonstrates that the recitation contest and web-scraped article datasets are more closely aligned. This similarity can be attributed to the fact that both datasets primarily consist of well-structured prose articles, which often explore topics related to local culture, customs and traditions. On the other hand, the content in Hokkien textbooks is composed of verses resembling children’s rhymes, making it more akin to Hokkien song lyrics. Subtitles exhibit a distinctive divergence as they primarily feature colloquial sentence structures, setting them apart from the other two groups.
A.2. JSD of Fine-tuning Parallel Data
In the parallel corpus, we adhere to the CKIP word segmentation system, which is designed for Traditional Chinese. Consequently, we calculate JSD similarity scores based on the Traditional Chinese portion of the text. Figure 4 presents the similarity between the training set, comprising dictionary and technical terms, and the test set, which includes iCorpus-100 and TAT. We observe that the technical terms subset diverges significantly from the others, because it consists solely of terminology, making it less similar to more general sentences. Moreover, there are significant domain differences between the training and test set, indicating that achieving a high performance on the translation with this test set presents a considerable challenge.
Appendix B Details of the GPT-4 evaluation methodology
B.1. Prompt Template
@@
B.2. Qualitative Analysis of GPT-4 Score Ranges and Examples
In the evaluation of translation scores from GPT-4, we performed a qualitative analysis of a sample of 20 examples from five distinct score intervals, as shown in Table 4. To offer a more comprehensive understanding, we have included examples of translations from Hokkien Han to Mandarin Chinese and English in Tables 8 and 9, respectively.
@@ B.2. Qualitative Analysis of GPT-4 Score Ranges and Examples
Hokkien Han Source: 準備離婚的翁仔姐仔共同攑槌,對準結婚手指殘殘摃落,佇觀眾的噗仔聲中,結婚手指當場歪斜走形。 English Reference: The couple who were preparing for divorce both lifted their hammers and struck hard at their wedding rings. Amidst the applause of the audience, the wedding rings twisted and deformed on the spot. |
| Model Translation Results |
GPT-4 Score |
@@
The couple prepared to divorce and took the hammer together. They hit each other’s wedding rings with force, causing them to deform in front of the audience’s applause. |
90 |
@@
The doll sisters prepared to divorce, and together held the hammer. They hit their wedding ring hard, causing it to bend in front of the audience’s applause. |
60 |
@@
The bride and groom held the hammer together to sign their divorce agreement. They both hit each other’s ring finger hard, causing it to twist in front of the audience’s applause. |
50 |
@@
The doll sisters who are preparing for divorce hold a hammer together, and the ring finger is broken in front of the audience’s applause. On stage, they immediately ran away with their hands crossed. |
20 |
@@
The bride and groom held the hammer together to hit the nails, but they accidentally hit each other’s fingers. In front of the audience, their hands were twisted in a strange way. |
10 |

### v2

@@ 5.1. Experimental Ablation Studies
In order to isolate the impact of various data inputs, we conducted ablation studies on continued pre-training of three different Hokkien writing systems, the extension of the input vocabulary with Hokkien Chinese characters and Hokkien Latin systems, and fine-tuning with these three different Hokkien writing systems141414Our focus on Hokkien LRL led us to exclude HRL data from training and evaluation..
@@ 5.1.1. Continued Pre-training Corpora and Vocabulary Extension Ablation Studies
As shown in Table 5, we took LLaMA 2-7B without any continued pre-training as a baseline. We compared its performance with the TAIDE-7B model continued pre-trained on different monolingual data and evaluated the impact of extending the Hokkien dictionary. Here, “NONE” indicates no continued pre-training on any Hokkien data, “HAN” indicates pre-training solely on Chinese character-based Hokkien data, and “ALL” indicates the inclusion of Latin script (POJ) and hybrid (HL) Hokkien data in addition to “HAN”. All these models were then fine-tuned using all available parallel data.
The baseline model, which has no further continued pre-training, performs the worst. In contrast, the TAIDE-7B model improves significantly in all translation directions without relying on Hokkien monolingual data but ZH data. This suggests that using a similar HRL model as a foundational model is beneficial when supplementary monolingual data is not available. Pre-training on HAN data improves the GPT-4 score by to points in HAN-related translations. Incorporating all Hokkien data yields the best performance, particularly in POJ-related translations, with a to points increase in GPT-4 score. These findings align with previous research (Yang et al., 2023; Li et al., 2023; Xu et al., 2023), demonstrating the substantial performance improvement in translation tasks when monolingual data is utilized for languages that the foundational model is not familiar with.
Regarding vocabulary extension, since the added vocabulary primarily consists of Latin scripts from POJ, the model with vocabulary extension exhibits superior performance only when the target language is POJ. For other translation directions, these models exhibit a slight decrement, averaging points lower on the GPT-4 score compared to models without vocabulary extension. We attribute this to the limited size of the pre-training corpus, which hinders effective tuning of newly added tokens. Consequently, we opted not to extend the vocabulary and suggest future work in collecting a larger POJ corpus for further investigation.
@@ 5.1.2. Fine-tuning Datasets Ablation Study
Table 6 indicates that models fine-tuned on parallel datasets significantly outperform few-shot ICL models across all translation directions, particularly in directions involving Latin script Hokkien (POJ), with a GPT-4 score increasing from to points. Demonstrate that the benefits of fine-tuning are particularly substantial for writing systems that are not closely related to HRL.
For ZHHAN and ENHAN directions, including ENHAN data enhances translation performance. However, further incorporation of POJ and HL parallel data does not yield additional improvements, suggesting that focusing solely on HRL in the parallel data is more effective for aligning cross-lingual embeddings.
In Hokkien script translation, the model performs the best when all parallel data are included. The inclusion of POJHL data dramatically improves translation capabilities in the POJHAN. The chrF++ improved by and points for HAN-POJ and POJ-HAN directions, respectively. We hypothesize that HL contains a few Han characters, allowing the model to learn some lexical correlations between different scripts in Hokkien.
@@ 5.2. Pre-training Corpus Script-Standardization
We evaluated the efficacy of continued pre-training using three different corpora:
1.
Monolingual Hokkien data in Chinese characters only.
2.
Monolingual Hokkien data in all writing systems.
3.
All monolingual Hokkien data standardized into Han characters.
These three models were subsequently fine-tuned using only ZHHAN and ENHAN parallel data.
Table 7 presents the evaluation results of these three models on TAT. The model pre-trained on standardized monolingual data on par or slightly outperforms the other models on average across both pairs of languages. This suggests that standardizing the writing system into Chinese characters can yield benefits on HAN translation.
@@ 6. Conclusion
We conducted a thorough examination of the efficiency of large language models in a translation system including Mandarin Chinese, English and two different Hokkien writing systems (HAN and POJ). We employed evaluation metrics including BLEU score, chrF++, and a GPT-4-based scoring method. Our results showed that employing Hokkien similar high-resource language model as a foundational model led to significant performance enhancements. Furthermore, models pre-trained on all available Hokkien data exhibited the highest performance. Extending vocabulary appeared to favor the incorporation of more monolingual data to elicit its efficiency. In the fine-tuning stage, the model demonstrates enhanced HAN translation only when it is supplied with parallel data corresponding to the HRLs that were extensively involved in its prior pre-training. Additionally, we investigated the benefits of standardizing the Hokkien scripts to Chinese characters. Our future work will explore the potential advantage of data augmentation through translations from ZH monolingual corpora into Hokkien Han and assess its impact on translation quality. Moreover, extending this research to include other prevalent spoken languages in Taiwan, like Hakka, can offer a more extensive viewpoint on handling the linguistic variations in Taiwan.
@@ 7. Limitation
The methodology used in this study is conditional on the fact that one writing system in a low-resource language is similar to a high-resource language, which can be seen as a limitation. In this work, Hokkien Han and Mandarin Chinese share similar writing systems and possess a substantial amount of common vocabulary. This allows for a transfer of knowledge from extensive Mandarin Chinese texts, thereby leveraging the benefits of large language models pre-trained on abundant Mandarin Chinese corpora to achieve an exceptionally efficient translation model.
@@ 9. Acknowledgements
We would like to express our sincere gratitude to Dr. Yu-Chun Wang for the generous assistance and insightful discussions on Taiwanese Hokkien. Our thanks also go to Shou-Yi, Hung for his invaluable help in polishing the paper writing. Additionally, we are deeply grateful to all the reviewers for their dedication to improving this research. We are thankful to the Trustworthy AI Dialogue Engine (TAIDE) project for providing the Traditional Chinese academic research foundation LLM, which served as a crucial base for building our model. Special appreciation goes to the National Center for High-performance Computing (NCHC) for providing computational and storage resources. This work was also supported in part by the Co-creation Platform of the Speech-AI Research Center, Industry-Academia Innovation School, NYCU, under the framework of the National Key Fields Industry-University Cooperation and Skilled Personnel Training Act, from the Ministry of Education (MOE), the National Development Fund (NDF), and industry partners in Taiwan.
@@ 10. Bibliographical References
- Brown et al. (2020) Tom Brown, Benjamin Mann, Nick Ryder, Melanie Subbiah, Jared D Kaplan, Prafulla Dhariwal, Arvind Neelakantan, Pranav Shyam, Girish Sastry, Amanda Askell, Sandhini Agarwal, Ariel Herbert-Voss, Gretchen Krueger, Tom Henighan, Rewon Child, Aditya Ramesh, Daniel Ziegler, Jeffrey Wu, Clemens Winter, Chris Hesse, Mark Chen, Eric Sigler, Mateusz Litwin, Scott Gray, Benjamin Chess, Jack Clark, Christopher Berner, Sam McCandlish, Alec Radford, Ilya Sutskever, and Dario Amodei. 2020. Language models are few-shot learners. In Advances in Neural Information Processing Systems, volume 33, pages 1877–1901. Curran Associates, Inc.
@@
- Rapp (2009) Reinhard Rapp. 2009. The backtranslation score: Automatic MT evalution at the sentence level without reference translations. In Proceedings of the ACL-IJCNLP 2009 Conference Short Papers, pages 133–136, Suntec, Singapore. Association for Computational Linguistics.
@@
Appendix A Details of the GPT-4 evaluation methodology
A.1. Prompt Template for Evaluations
@@
A.2. Examples of GPT-4 Scoring on Translations
To provide a more comprehensive understanding of the GPT-4 scoring standards, we have included examples of translations from Hokkien Han to Mandarin Chinese and English in Table 8 and Table 9, respectively. These examples demonstrate the varied scoring outcomes provided by GPT-4, alongside the translation qualities that correspond to different scores.
@@ A.2. Examples of GPT-4 Scoring on Translations
| Hokkien Han Source: 準備離婚的翁仔姐仔共同攑槌,對準結婚手指殘殘摃落,佇觀眾的噗仔聲中,結婚手指當場歪斜走形。 English Reference: The couple who were preparing for divorce both lifted their hammers and struck hard at their wedding rings. Amidst the applause of the audience, the wedding rings twisted and deformed on the spot. | |
| Model Translation Results | GPT-4 Score |
@@
| The couple prepared to divorce and took the hammer together. They hit each other’s wedding rings with force, causing them to deform in front of the audience’s applause. | 90 |
@@
| The doll sisters prepared to divorce, and together held the hammer. They hit their wedding ring hard, causing it to bend in front of the audience’s applause. | 60 |
@@
| The bride and groom held the hammer together to sign their divorce agreement. They both hit each other’s ring finger hard, causing it to twist in front of the audience’s applause. | 50 |
@@
| The doll sisters who are preparing for divorce hold a hammer together, and the ring finger is broken in front of the audience’s applause. On stage, they immediately ran away with their hands crossed. | 20 |
@@
| The bride and groom held the hammer together to hit the nails, but they accidentally hit each other’s fingers. In front of the audience, their hands were twisted in a strange way. | 10 |
Appendix B JSD for Corpora
B.1. JSD of Continued Pre-training Monolingual Corpora
We use Jensen-Shannon divergence (JSD) as a metric to assess the domain similarity of the monolingual HAN corpora. Due to the absence of open-sourced word segmentation tools for HAN, calculating JSD presents a notable challenge. However, the HAN writing system’s resemblance to Traditional Chinese characters allows us to utilize the CKIP Traditional Chinese segmentation tool151515https://github.com/ckiplab/ckiptagger for processing the HAN corpora. We then computed the JSD and analyzed the domain similarity of our corpora.
Figure 3 demonstrates that the recitation contest and web-scraped article corpora are more closely aligned. This similarity can be attributed to the fact that both corpora primarily consist of well-structured prose articles, which often explore topics related to local culture, customs and traditions. The content in Hokkien textbooks is composed of verses resembling children’s rhymes, making it more akin to Hokkien song lyrics. Subtitles exhibit a distinctive divergence as they primarily feature colloquial sentence structures, setting them apart from the other two groups.
B.2. JSD of Fine-tuning Parallel Data
In the parallel datasets, the CKIP word segmentation tool is employed to process the Traditional Chinese texts. Consequently, we calculate JSD similarity scores based on the Traditional Chinese portion of the texts. Figure 4 illustrates the similarity between the training set, which includes both dictionary and technical terms, and the test set, encompassing iCorpus-100 and TAT. We observe that the technical terms subset diverges significantly from the others, because it consists solely of terminology, making it less similar to more general sentences. Moreover, there are significant domain differences between the training and test sets, indicating that achieving a high performance on the translation with this test set presents a considerable challenge.
Appendix C Evaluating GPT-4’s Translation Performance on Hokkien
To evaluate GPT-4’s translation performance on Hokkien, we conducted experiments prompting it to translate between Hokkien and both ZH and English. When prompted to translate into HAN, GPT-4 predominantly generated output in ZH, with a limited mixture of HAN and Cantonese words. This finding led us to conclude that this approach is not suitable for assessing back-translation accuracy, as it primarily evaluates GPT-4’s translation capabilities in ZH, rather than Hokkien. Additionally, when prompted to translate into POJ, GPT-4’s output was completely incomprehensible.
Consequently, we only present the results where the target language is ZH or English in Table 10. When translating from HAN, GPT-4 outperforms our best model by and points on the GPT-4 score for HAN-ZH and HAN-EN translation tasks, respectively. Apart from its significantly larger model size, GPT-4’s superior performance might be attributed to the similarity writing system between ZH and HAN161616When directly comparing HAN and ZH sentences in iCorpus-100 dataset, we obtain a BLEU score of and chrF++ of ., allowing it to process HAN as a noisy version of ZH and leverage its knowledge of ZH. In contrast, when the source language is POJ, GPT-4 struggles to produce meaningful translations, performing worse than our model with GPT-4 scores of and points for POJ-ZH and POJ-EN, respectively. This emphasizes the need for a specialized large language model designed for Hokkien, which this research aims to address.
| Model | BLEU | chrF++ |
| HAN-ZH | ||||||||
| Our Best | 32.60 | 34.91 | 75.35 | 63 | ||||
| GPT-4 | 59.47 | 59.07 | 86.50 | 81 | ||||
| HAN-EN | ||||||||
| Our Best | 18.45 | 44.92 | 66.75 | 38 | ||||
| GPT-4 | 33.74 | 58.36 | 82.70 | 79 | ||||
| POJ-ZH | ||||||||
| Our Best | 30.08 | 31.36 | 43.10 | 12 | ||||
| GPT-4 | 12.19 | 16.65 | 5.45 | 1 | ||||
| POJ-EN | ||||||||
| Our Best | 7.83 | 30.53 | 24.20 | 3 | ||||
| GPT-4 | 1.78 | 22.69 | 3.80 | 0 | ||||